OpenStudy (anonymous):

-3√64 find real roots

OpenStudy (callisto):

\[\sqrt{64}=\sqrt{8^2} =8\]Therefore\[\-3sqrt{64}=-3\sqrt{8^2} =-3(8)\]Can you work it out?

OpenStudy (anonymous):

-24 and 24

OpenStudy (anonymous):

isn't it .25? a website calculator gave me that answer..

OpenStudy (callisto):

Revised: \[\sqrt{64} = \pm8\]Therefore\[-3\sqrt{64} =-3( \pm8)\] = -3(8) or -3 (-8) = -24 or 24 Sorry my mistakes!

OpenStudy (anonymous):

its not -3 times the square root of 64 its negative cube root of 64

OpenStudy (anonymous):

\[\huge \sqrt[-3]{64}\] like this??

OpenStudy (anonymous):

yes, my bad i made it look like -3 times the square root 64 lol

OpenStudy (callisto):

\[-\sqrt[3]{64}= -\sqrt[3]{4^3} = -4\]

OpenStudy (anonymous):

Thanks!!!

OpenStudy (callisto):

Welcome, but do you understand?

OpenStudy (anonymous):

yes :] i just got confused lol

OpenStudy (anonymous):

Hold on... is it \[\huge \sqrt[-3]{64} \quad or \quad -\sqrt[3]{64}\]

OpenStudy (anonymous):

first one

OpenStudy (anonymous):

how do u make it look like that?

OpenStudy (callisto):

\[\sqrt[-3]{64} = 1/\sqrt[3]{64} = 1/4\] My bad :(

OpenStudy (anonymous):

wait wtf.

OpenStudy (anonymous):

ugh now im confused again =[ i thought it was -4 and the website amde a mistake =/

OpenStudy (callisto):

Sorry i misunderstood your question twice..

OpenStudy (anonymous):

made*

OpenStudy (anonymous):

\[\huge \sqrt[-3]{64}=\sqrt[-3]{2^8} =\left(2\right)^{\frac{8}{-3}}\]

OpenStudy (anonymous):

64 cubed is 4.. why did u put it s 2^8?

OpenStudy (anonymous):

I screw them up always ! LOL ... now I see it ! :S

OpenStudy (anonymous):

lol its kewl, where the flutter is callisto??

OpenStudy (anonymous):

\[\huge \sqrt[-3]{64}=\left(4\right)^{\frac{3}{-3}}=4^{-1}=\frac14\]

OpenStudy (anonymous):

Thanks bro I get it now!!

OpenStudy (anonymous):

Glad to help :)

OpenStudy (callisto):

my English fails me :(

OpenStudy (anonymous):

@Callisto You're not British\American ??

OpenStudy (callisto):

Neither of them :(

OpenStudy (anonymous):

where do you live?

OpenStudy (anonymous):

hey just a quick question can there be a negetive number in the radicand of a cube root?

OpenStudy (anonymous):

draw it.. I don't understand it at all !